Problem1 Suppose we have an econonny that produces ony oge and coconnts. The following table gives the prices and qasntities of these commoditis for 2017 and 2018 2017 2018 Oranges 13 2 of 5 Price 110 Quantity Value Coconuts Price Quantity Value Nominal GDP 100 1 Fill-in the missing values in the table. 1.2 Caleulate the gnowth rate in sominal GDP etwees 2017 and 2018 Growth rate 13 Compute the value of GDP for 2018 using 2017 peis Then compute the growth rate in real GDP from 2017 to 2018 based on 2017 prices 2018 GDP (at 2017 prices) 2017 GDP (at 2017 prices) Growth rate 1.4 Congaste t be value of GDP fr 2017 tsing 2018 pekrs. T in real GDP from 2017 to 2018 based on 2018 prions n compute t노 growth rate 2017 GDP (at 2018 prices) 2018 GDP (at 2018 prices) Growth rate 15 What do you notice about the resalts from using the different base year prices? D they give the sane increase in real GDP Answer
Answer #1

Answer :1.1)  For 2017 :

Value of Oranges = $ (1×100) = $100

Value of Coconuts = $ (3×20) = $60

Nominal GDP = $100 + $60 = $160.

For 2018 :

Value of Oranges = $ (1.5 × 110) = $165

Value of Coconuts = $ (2 × 50) = $100

Nominal GDP = $165 + $100 = $265.

1.2) Growth Rate = [(Nominal GDP of 2018 - Nominal GDP of 2017) / Nominal GDP of 2017] × 100

=> Growth Rate = [(265 - 160) / 160] × 100 = 65.6%

Therefore, Growth Rate = 65.6%.

1.3) 2018 GDP (at 2017 price) = (Price of 2017 × Quantity of 2018 for Oranges) + (Price of 2017 × Quantity of 2018 for Coconuts)

=> 2018 GDP = (1 × 110) + (3 × 50) = 110 + 150

Therefore, 2018 GDP (at 2017 price) = 260.

2017 GDP (at 2017 price) = (1 × 100) + (3 × 20) = $160.

Growth Rate = [(260 - 160) / 160] × 100

Therefore, Growth Rate = 62.5%

1.4) 2017 GDP (at 2018 price) = (1.5 × 100) + (2 × 20) = 150 + 40 = 190

Therefore, 2017 GDP (at 2018 price) = 190.

2018 GDP (at 2018 price) = (1.5 × 110) + (2 × 50) = 265

Therefore, 2018 GDP (at 2018 price) = 265.

Growth Rate = [(265 - 190) / 190] × 100 = 39.5%

Therefore, Growth Rate = 39.5%

1.5) The difference in result at different base year prices is that the growth rate at 2017 price level is higher than the growth rate at 2018 price level.

No. There is no same increase in Real GDP level. Because the growth rate at 2017 price is 62.5% and the growth rate at 2018 is 39.5%.